OpenStudy (anonymous):

taylor series of f(x)=ln(x), a=2

OpenStudy (ash2326):

Taylor's Series expansion of f(x) about the point a \[f(x)=f(a)+f'(a) (x-a)+\frac{f''(a)(x-a)^2}{2!}+\frac{f'''(a)(x-a)^3}{3!}....\] Can you do this now?

OpenStudy (anonymous):

Yes, my first intuition was to do this, but I thought that I was doing "too much work"? How would I go about finding the radius of convergence? Thank you!

OpenStudy (ash2326):

@caaaitlin I did some googling and found this page. It's helpful. Please go through this and if you have any doubt do tell me http://www.sosmath.com/calculus/tayser/tayser01/tayser01.html